OpenStudy (anonymous):

Which one is a complex number in standard form for the expression 2-5i/1+9i?

OpenStudy (anonymous):

z=2-5i/1+9i z=2-5i/1+9i*1-9i/1-9i z=2-18i-5i+45i^2/1-81i^2 z=2-23i+45*-1/1-81*-1 z=2-23i-45/1+81 z=-23i-43/82

OpenStudy (anonymous):

thankyou